OpenStudy (anonymous):

Sharmi and Reese each deposit money into a savings account. Sharmi deposits $200 less than Reese. After the same amount of time, Reese has earned $90 in interest at a rate of 3% per year, and Sharmi has earned $100 in interest at a rate of 5% per year. How much money did Reese and Sharmi initially deposit into their savings accounts?

OpenStudy (anonymous):

flutter YOU sleek-feathered one pellet wingspan HEAD

OpenStudy (anonymous):

HURRY UP YOU PIECE OF pellet!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

BURN IN HELL

OpenStudy (anonymous):

F U C K

OpenStudy (anonymous):

YOUR FAMILY LIVES OFF OF WELFARE

OpenStudy (anonymous):

TOOL AS retrice HOLE MOTHER flutterER wingspan

OpenStudy (mathstudent55):

Let the amount Reese deposited be x Sharmi's deposit was x - 200 Interest = Prt (P = principal amount, r = rate, t = time) Reese earned $90 in interest, so Prt = x * 0.03 * t = 90 Sharmi earned $100 in interest, so Prt = (x - 200) * 0.05 * t = 100 You have two equations and two unknowns, so you can solve it as a system of equations. x * 0.03 * t = 90 (x - 200) * 0.05 * t = 100